# Continuous Language Model Interpolation for Dynamic and Controllable Text Generation

### v1

@@ Abstract
As large language models (LLMs) have gained popularity for a variety of use cases, making them adaptable and controllable has become increasingly important, especially for user-facing applications. While the existing literature on LLM adaptation primarily focuses on finding a model (or models) that optimizes a single predefined objective, here we focus on the challenging case where the model must dynamically adapt to diverse —and often changing— user preferences. For this, we leverage adaptation methods based on linear weight interpolation, casting them as continuous multi-domain interpolators that produce models with specific prescribed generation characteristics on-the-fly. Specifically, we use low-rank updates to fine-tune a base model to various different domains, yielding a set of anchor models with distinct generation profiles. Then, we use the weight updates of these anchor models to parametrize the entire (infinite) class of models contained within their convex hull. We empirically show that varying the interpolation weights yields predictable and consistent change in the model outputs with respect to all of the controlled attributes. We find that there is little entanglement between most attributes and identify and discuss the pairs of attributes for which this is not the case. Our results suggest that linearly interpolating between the weights of fine-tuned models facilitates predictable, fine-grained control of model outputs with respect to multiple stylistic characteristics simultaneously. [^1] Code: https://github.com/skangasl/continuous-lm-interpolation
@@ 1 Introduction
Large language models (LLMs) are used for a diverse set of applications due to their high performance across a wide spectrum of tasks (Bubeck et al., 2023). In many common LLM use cases (such as chatbots), different users often have distinct and continuously evolving preferences for the type of output they want. For example, a user might want a creative and verbose response for certain queries, but a concise and precise response for others. In practice, a user may try different variations of the same query succesively until they elicit a generation that matches their goal. This trial-and-error process can be time-consuming and lacks guaranteed results, especially since minor word changes in a prompt can have disproportionate impact on the output. Additionally, expressing fine-grained continuous preferences (e.g., simplicity of the response) is often difficult in —inherently discrete— natural language. These challenges are exacerbated when the user has complex, multi-faceted preferences (e.g., a specific combination of simplicity, formality, and verbosity) that they expect the generation to satisfy all at once. As a result, there is a pressing need for methods that allow for fine-grained and predictable control over LLM text generation, and which can adapt on-the-fly to mutable user preferences and constraints.
@@
Prior work in controllable text generation (CTG) has largely focused on optimizing for one set of control criteria through techniques such as instruction tuning (Zhou et al., 2023), modifying the output probability distributions (Pascual et al., 2021; Yang & Klein, 2021; Dekoninck et al., 2024), changing model activations at inference time (Li et al., 2023), learning modifications to the embeddings (Li & Liang, 2021; Han et al., 2023), or training (Keskar et al., 2019; Krause et al., 2021). These methods, however, do not naturally allow for the composition of multiple objectives and lack fine-grained control, especially those that rely on the user expressing preferences in natural language, for the reasons described above. Embedding modification and inference-time approaches do not allow for as complex tuning to the objective as fine-tuning based ones and often require additional training for each control variable value. While fine-tuning to each desired objective would likely allow for the most precise optimization, this is computationally infeasible to do for each combination of control variables and strengths of control in the entire (infinite) set of possible combinations.
@@
With these challenges in mind, here we seek to enable dynamic and controllable text generation in a manner that takes advantage of the strengths of fine-tuning while remaining computationally feasible for dynamically changing control variables. Recent work has demonstrated that multiple pre-trained or fine-tuned models can be effectively composed through linear weight interpolation (Wortsman et al., 2022; Ilharco et al., 2023). This has also been shown to extend to models trained with parameter-efficient fine-tuning (PEFT) methods (Zhang et al., 2023; Huang et al., 2024) such as low-rank adaptation (Hu et al., 2021). We build upon and extend this line of work by showing that linear weight interpolation can be used to obtain models with specific mixtures of characteristics on-the-fly and without additional training, effectively providing a continuous parametrization of the (infinite) ‘convex hull’ of a set of fine-tuned models. To do so, we fine-tune two endpoint anchor models for each control attribute, one at each extreme of attribute strength. We then interpolate along the vector between the weights of these two models for each attribute before computing a weighted average across all of the single-attribute interpolated models. Thus, varying the interpolation and averaging weights gives us dense coverage of the model parameter space, allowing us to create models tailored to any preference profile spanned by the fine-tuned models. We evaluate linear weight interpolation for multiple style attributes and demonstrate empirically that changes in the interpolation and averaging weights yield predictable and consistent responses in the level of each attribute in the generated text.
Figure 1: Overview of our continuous weight interpolation formulation. The blue boxes represent the inputs controlled on-the-fly by the user. When the user chooses $\alpha$ and $\lambda$, we first interpolate between the weights $\theta_{+i}$ and $\theta_{-i}$ of two fine-tuned endpoint models for each control $i$ using $\alpha_{i}$, then weight each interpolated model $\theta_{\alpha_{i}}$ by $\lambda_{i}$ and sum over all $\lambda_{i}\theta_{\alpha_{i}}$. In this manner, the user dynamically controls attributes of the output by specifying $\alpha$ and $\lambda$.
@@
A potential pitfall of this approach is that, as seen in prior work in the vision domain (Ortiz-Jimenez et al., 2023), the weights for different single-attribute interpolated models may be entangled. This could lead to unexpected correlations between attributes in the averaged models. These correlations are detrimental to CTG, as changing the interpolation weights for one attribute could have an unexpected effect on the correlated attributes in the output text. However, we find that there is surprisingly little entanglement between the vast majority of control attributes and analyze the pairs of controls where this is not the case.
@@
In summary, our key contributions are: (1) we show how parameter-efficient adaptation methods can be used to continuously interpolate between models fine-tuned with various distinct generation objectives, allowing for on-the-fly adaptation to user-specified generation preferences expressed in terms of interpretable control variables; and (2) we demonstrate that changes in the interpolation yield smooth and predictable changes in the properties of the generated text across multiple sets of controls with limited entanglement.
@@ 2.1 Datasets
For each style attribute, we fine-tune a separate anchor Llama2-7b model (Touvron et al., 2023) on two datasets representing the extremes of the attribute level. For simplicity, we use the TinyStories dataset (Eldan & Li, 2023) to fine-tune a simple model and novel chapters from the BookSum dataset (Kryscinski et al., 2021) to fine-tune a complex model. We use the documents classified as formal and informal in Grammarly’s Yahoo Answers Formality Corpus (GYAFC) dataset (Rao & Tetreault, 2018) to fine-tune formal and informal models. For the politeness attribute, we use the documents in the highest and lowest politeness class in the work by Madaan et al. (2020) for fine-tuning polite and impolite models, respectively. We fine-tune positive and negative sentiment models using the Stanford Sentiment Treebank (SST-2) dataset (Socher et al., 2013). For humor, we use the FlickrStyle dataset (Gan et al., 2017) to fine-tune humorous and non-humorous models.
@@ 2.2 Fine-tuning
We employ Low-Rank Adaptation (LoRA) in order to fine-tune our models in a parameter-efficient manner (Hu et al., 2021). In LoRA fine-tuning, at each layer of the transformer model, the pretrained model weights are frozen and low-rank decomposition matrices are learned to adapt the model in fine-tuning. Denoting the pretrained language model weights as $\theta_{PRE}\in\mathbb{R}^{d_{1}\times d_{1}}$, LoRA computes the updated weights as follows:
@@
Here, $A\in\mathbb{R}^{k\times d_{2}}$ and $B\in\mathbb{R}^{d_{1}\times k}$ (with $k\ll d_{1},d_{2}$) are trainable parameters learned during fine-tuning. We use LoRA as an adaptation method because it requires significantly fewer parameters than traditional fine-tuning while maintaining similar performance, so LoRA weights can be quickly modified and applied to large pretrained language models. We use the parameters in Appendix A.1 for fine-tuning the models and fine-tune two LoRA models per style characteristic, one on each of the extreme classes outlined in 2.1.
@@ 2.3 Linear weight interpolation
We formulate linear weight interpolation between the LoRA fine-tuned models in terms of interpolation weights $\alpha_{i}$ and attribute mixing weights $\lambda_{i}$ as shown in Figure 1. We denote $\theta_{+i}=\theta_{PRE}+B_{+i}A_{+1}$ and $\theta_{-i}=\theta_{PRE}+B_{-i}A_{-1}$ as the two LoRA fine-tuned endpoint anchor models for attribute $i$. Then, for a single attribute, we interpolate along the vector between the two fine-tuned endpoint models by computing
$\theta_{\alpha_{i}}$ $=\theta_{PRE}+\alpha_{i}\theta_{+i}+(1-\alpha_{i})\theta_{-i}=\theta_{PRE}+\alpha_{i}B_{+i}A_{+i}+(1-\alpha_{i})B_{-i}A_{-i}$ (2)
@@
$\theta_{\alpha,\lambda}$ $=\sum_{i}\lambda_{i}\theta_{\alpha_{i}}=\theta_{PRE}+\sum_{i}\lambda_{i}(\alpha_{i}B_{+i}A_{+i}+(1-\alpha_{i})B_{-i}A_{-i})$ (3)
We denote $\lambda_{i}$ to be the mixing weight for the $i$th attribute and constrain $\sum_{i}\lambda_{i}=1$. We note that the case with one attribute dimension corresponds to the sum having a single term with $\lambda_{1}=1$. With this formulation, we can construct any model in the convex hull of the fine-tuned models by choosing appropriate interpolation weights $\alpha$ and mixing weights $\lambda$.
(a) Interpolation between the anchor models
(b) Linear extrapolation beyond the anchor models
Figure 2: Effect of linear weight interpolation for a single attribute dimension. For each style attribute, we report the attribute score when linearly interpolating between the models with weight $\alpha$. We show the effect of 1(a) interpolating between the two fine-tuned models and 1(b) extrapolating beyond the fine-tuned models ($\alpha<0$ and $\alpha>1$). Increasing $\alpha$ between the two fine-tuned models results in a smooth increase in the attribute score. When extrapolating, there is a stable region where this trend continues until a certain point (around $\alpha$ equal to $-1$ and $2$), where performance degrades and the extrapolation is unstable.
Figure 3: Wikitext perplexity of linearly interpolated and extrapolated models. We report the average perplexity (lower is better) of each model from Figure 1(b) on the Wikitext test set. For all of the interpolated models ($\alpha\in[0,1]$), the perplexity is either better than or between the performance of the two fine-tuned models. For the extrapolated models ($\alpha<0.0$ and $\alpha>1.0$), the perplexity increases rapidly beyond $\alpha$ values of around $-1$ and $2$. We clip the y-axis at 7.0 for readability (the full plot is shown in Figure 9).
@@ 2.4 Evaluation
To evaluate the generations of each interpolated model, we use a subset of 1k randomly sampled prompts from the WritingPrompts dataset (Fan et al., 2018) and generate 3 continuations for each prompt. We compute scores for each of the attributes to evaluate the level of each control criterion. Similarly to prior work on text style transfer (Xu et al., 2018), we fine-tune a RoBERTa (Liu et al., 2019) classification head on each attribute and compute a sigmoid over the output logits to obtain the probability of class $1$, which we report as the attribute score. We label the documents such that an attribute score closer to $1$ corresponds to a document that is more simple, formal, polite, positive in sentiment, or humorous. We also compute perplexity on the test split of the WikiText dataset (Merity et al., 2016) to evaluate model fluency.
@@ 3 Continuous Language Model Interpolation
We begin by investigating the linear interpolations between each pair of low-rank fine-tuned anchor models (3.1). We then extend this analysis to the convex hull of fine-tuned models for multiple attributes (3.2).
Figure 4: Effect of $\alpha_{i}$ and $\lambda_{i}$ on attribute scores for multiple control dimensions. In each plot, we report the attribute scores for multiple $\alpha_{i}$ values when $\alpha_{i}$ is varied for the attribute $i$ in the column title and kept constant at $0.0$ for the remaining attributes. $\lambda_{i}$ for the control dimension $i$ for which the $\alpha_{i}$ weight is being varied is listed in the row titles. The mixture weights $\lambda_{j}$ for all remaining $j$ are set uniformly so that the sum of $\lambda$ weights is $1$ ($\lambda_{j}=(1.0-\lambda_{i})/4$). For all of the attribute dimensions, increasing $\alpha_{i}$ results in an increasing attribute score for that control dimension with limited effect on the other attribute scores.

### v2

# Continuous Language Model Interpolation yields Dynamic and Controllable Text Generation

@@ Abstract
As large language models (LLMs) have gained popularity for a variety of use cases, making them adaptable and controllable has become increasingly important, especially for user-facing applications. In particular, linear interpolation between model parameters forms the backbone for many recent approaches to adapting models to user preferences. While the existing literature on LLM adaptation primarily focuses on finding methods that optimize for some set of performance criteria or user preferences, here we instead seek to better understand and characterize the behavior of dense, continuous interpolation between models. Specifically, we use low-rank updates to fine-tune a base model to various different domains, yielding a set of anchor models with distinct generation profiles. Then, we use the weight updates of these anchor models to parametrize the entire (infinite) class of models contained within their convex hull. We empirically show that varying the interpolation weights yields predictable and consistent change in the model outputs with respect to all of the controlled attributes simultaneously. We find that there is little entanglement between most attributes and identify and discuss the pairs of attributes for which this is not the case. Our results suggest that parameter merging facilitates flexible model adaptation due to its predictable behavior within the full interpolation region. [^1] Code: https://github.com/skangasl/continuous-lm-interpolation
@@ 1 Introduction
Large language models (LLMs) are used for a diverse set of applications due to their high performance across a wide spectrum of tasks (Bubeck et al., 2023). In many common LLM use cases (such as chatbots), different users often have distinct and continuously evolving preferences for the type of output they want. For example, a user might want a creative and verbose response for certain queries, but a concise and precise response for others. In practice, a user may try different variations of the same query successively until they elicit a desired generation. This trial-and-error process can be time-consuming and lacks guaranteed results, especially since minor word changes in a prompt can have disproportionate impact on the output. Additionally, expressing fine-grained continuous preferences (e.g., simplifying a response by 25%) is often difficult in —inherently discrete— natural language. These challenges are exacerbated when the user has complex, multi-faceted preferences (e.g., a specific combination of simplicity, formality, and verbosity) that they expect the generation to satisfy all at once. As a result, it is critical to understand how to adapt LLMs to user preferences and constraints in a fine-grained and predictable way.
Figure 1: Overview of our continuous model interpolation framework. Given a collection of ‘anchor’ models fine-tuned on datasets at opposite ends of an attribute spectrum (e.g., $\theta_{+i}$: positive and $\theta_{-i}$: negative sentiment) for $N$ different attributes, the user selects interpolation parameters $\alpha$ (per-attribute spectrum modulation) and $\lambda$ (attribute mixture weights), which are used to generate a model with weights $\theta_{\alpha,\lambda}$ tailored to that specific parameter choice. We use this framework to analyze the behavior of interpolated models within the entire continuous region between the anchor models.
@@
Prior work in controllable text generation (CTG) has largely focused on optimizing for one set of control criteria through techniques such as instruction tuning (Zhou et al., 2023), modifying the output probability distributions (Pascual et al., 2021; Yang & Klein, 2021; Dekoninck et al., 2024), changing model activations at inference time (Li et al., 2023), learning modifications to the embeddings (Li & Liang, 2021; Han et al., 2023), or training (Keskar et al., 2019; Krause et al., 2021). These methods, however, are not parametrized continuously and instead require a fixed set of controls criteria. Thus, to achieve fine-grained control in the range between different attribute classes, they would have to be individually run for each specific set of intermediate attribute values, which is prohibitively expensive over a continuous range. Similarly, fine-tuning models with data that contains a proportionate amount of documents from each desired objective (ie $0.5$ positive and $0.5$ negative sentiment documents for a neutral model) would allow for the most precise optimization. However, this is computationally infeasible to do for each combination of control variables and strengths of control in the entire (infinite) set of possible combinations.
With these challenges in mind, linear weight interpolation has been proposed to adapt LLMs in a manner that takes advantage of the strengths of fine-tuning while making it computationally feasible to dynamically adapt the model for specific tasks or user preferences. Recent work has demonstrated that multiple pre-trained or fine-tuned models can be effectively composed through linear weight interpolation (Wortsman et al., 2022; Ilharco et al., 2023). This has also been shown to extend to models trained with parameter-efficient fine-tuning (PEFT) methods (Zhang et al., 2023; Huang et al., 2024) such as low-rank adaptation (Hu et al., 2021). As a result, linear interpolation has been used to adapt models through improving multitask performance (Matena & Raffel, 2021; Yadav et al., 2023; Ortiz-Jimenez et al., 2023) or aligning models to user preferences in the reinforcement learning setting (Ramé et al., 2023; Jang et al., 2023; Wang et al., 2024). However, these approaches largely center on optimizing the interpolation for task performance or user reward, so the behavior of fine-grained interpolation in the full continuous region between models remains poorly understood.
@@
In this work, we show that linear weight interpolation effectively provides a continuous parametrization of the (infinite) ‘convex hull’ of a set of fine-tuned models. To do so, we fine-tune two endpoint anchor models for each control attribute, one at each extreme of attribute strength. We then interpolate along the vector between the weights of these two models for each attribute before computing a weighted average across all of the single-attribute interpolated models (Figure 1). Thus, varying the interpolation and averaging weights gives us dense coverage of the parameter space between endpoint models. We evaluate linear weight interpolation for multiple style attributes and demonstrate empirically that changes in the interpolation and averaging weights yield predictable and consistent responses in each control attribute in the generations.
@@
A potential pitfall of linear interpolation is that, as seen in prior work in the vision domain (Ortiz-Jimenez et al., 2023), the weights for different single-attribute interpolated models may be entangled. This could lead to unexpected correlations between attributes in the averaged models. These correlations are detrimental to controllability, as changing the interpolation weights for one attribute could have an unexpected effect on the correlated attributes in the output text. However, we find that there is surprisingly little entanglement between the vast majority of control attributes and analyze the pairs of controls where this is not the case.
@@
In summary, our key contributions are: (1) we provide a framework for analyzing parameter-efficient adaptation in the continuous interpolation region between models fine-tuned with various distinct generation objectives; and (2) we demonstrate that changes in the interpolation yield smooth and predictable changes in the properties of the generated text across multiple sets of controls with limited entanglement.
@@ 2.1 Datasets
For each style attribute, we fine-tune a separate anchor Llama2-7b model (Touvron et al., 2023) on two English datasets representing the extremes of the attribute level. For simplicity, we use the TinyStories dataset (Eldan & Li, 2023) to fine-tune a simple model and novel chapters from the BookSum dataset (Kryscinski et al., 2021) to fine-tune a complex model. We use the documents classified as formal and informal in Grammarly’s Yahoo Answers Formality Corpus (GYAFC) dataset (Rao & Tetreault, 2018) to fine-tune formal and informal models. For the politeness attribute, we use the documents in the highest and lowest politeness class in the work by Madaan et al. (2020) for fine-tuning polite and impolite models, respectively. We fine-tune positive and negative sentiment models using the Stanford Sentiment Treebank (SST-2) dataset (Socher et al., 2013). For humor, we use the FlickrStyle dataset (Gan et al., 2017) to fine-tune humorous and non-humorous models.
@@ 2.2 Fine-tuning
We fine-tune our models in a parameter-efficient manner using Low-Rank Adaptation (LoRA, Hu et al., 2021), which keeps pretrained model weights frozen but learns an additive low-rank matrix update for each layer during fine-tuning. Denoting the pretrained language model weights as $\theta_{PRE}\in\mathbb{R}^{d_{1}\times d_{2}}$, LoRA computes the updated weights as:
@@
Here, $A\in\mathbb{R}^{k\times d_{2}}$ and $B\in\mathbb{R}^{d_{1}\times k}$ (with $k\ll d_{1},d_{2}$) are trainable parameters learned during fine-tuning. We use LoRA as an adaptation method because it requires significantly fewer parameters than traditional fine-tuning while maintaining similar performance, so LoRA weights can be quickly modified and applied to large pretrained language models. We use the parameters in Appendix A.1 for fine-tuning the models and fine-tune two LoRA models per style characteristic, one on each of the extreme classes outlined in 2.1. We denote the two LoRA fine-tuned endpoint anchor models for attribute $i$ by $\theta_{+i}=\theta_{PRE}+B_{+i}A_{+i}$ and $\theta_{-i}=\theta_{PRE}+B_{-i}A_{-i}$.
@@ 2.3 Linear weight interpolation
Given a collection of fine-tuned model weights obtained by LoRA as described above, we generate interpolated models by linearly interpolating between their weights. We formulate linear weight interpolation between the LoRA fine-tuned models in terms of interpolation weights $\alpha_{i}$ and attribute mixing weights $\lambda_{i}$ as shown in Figure 1. For a single attribute, we interpolate along the vector between the two fine-tuned endpoint models by computing
$\theta_{\alpha_{i}}$ $=\alpha_{i}\theta_{+i}+(1-\alpha_{i})\theta_{-i}$ (2)
$=\theta_{PRE}+\alpha_{i}B_{+i}A_{+i}+(1-\alpha_{i})B_{-i}A_{-i}$
@@
$\theta_{\alpha,\lambda}$ $=\sum_{i}\lambda_{i}\theta_{\alpha_{i}}$ (3)
We denote $\lambda_{i}$ to be the mixing weight for the $i$th attribute and constrain $\sum_{i}\lambda_{i}=1$. We note that the case with one attribute dimension corresponds to the sum having a single term with $\lambda_{1}=1$. With this formulation, we can construct any model in the convex hull of the fine-tuned models by choosing appropriate interpolation weights $\alpha$ and mixing weights $\lambda$. While the raw interpolation parameters do not have a clear meaning, we seek to show that a user can controllably increase or decrease the level of each attribute by changing $\alpha$ and $\lambda$.
@@ 2.4 Evaluation
To evaluate the interpolated models, we use a subset of 1k randomly sampled prompts from the WritingPrompts dataset (Fan et al., 2018) and generate 3 continuations for each prompt. We compute scores for each attribute to evaluate the level of the control criterion. Similarly to prior work on text style transfer (Xu et al., 2018), we fine-tune a RoBERTa (Liu et al., 2019) classification head on each attribute using a held out split of the datasets in 2.1 and compute a sigmoid over the output logits to obtain the probability of class $1$, which we report as the attribute score. We label the documents such that an attribute score closer to $1$ corresponds to a document that is more simple, formal, polite, positive in sentiment, or humorous. We also compute perplexity on the test split of the WikiText dataset (Merity et al., 2016) and n-gram diversity of the WritingPrompts generations to evaluate text quality.
Baselines: We provide a comparison of weight interpolation to two baselines: DExperts (Liu et al., 2021) and model arithmetic (Dekoninck et al., 2024), as these are the main prior approaches that do not use linear weight interpolation but allow for continuous values of attribute strength parameter. DExperts is formulated using a base model and two fine-tuned endpoint models. Then, given a prompt $x_{<t}$, if we denote the output logits of the base model at time $t$ as $z_{t}$ and the logits of the two endpoint models as $z_{t}^{+}$ and $z_{t}^{-}$, respectively, then the DExperts output probability distribution is defined by $P(X_{t}|x_{<t})=\text{softmax}(z_{t}+\alpha(z_{t}^{+}-z_{t}^{-}))$. The model arithmetic baseline uses the same formula for computing the output probability distribution, but instead of using two fine-tuned endpoint models, it uses two prompt-conditioned endpoint models to produce $z_{t}^{+}$ and $z_{t}^{-}$. As a result, both DExperts and weight interpolation require $2*\text{number dimensions}$ fine-tuned endpoint models, while model arithmetic does not require any. However, weight interpolation uses a single inference pass, while both of these approaches require $2*\text{number dimensions}+1$ inference passes (Table 4).
We use the fine-tuned anchor models for DExperts and prompt-condition Llama-2-7b (see A.3 for details) for model arithmetic. For both comparisons, we compute the attribute scores for $\alpha\in[-2,2]$. To provide a direct comparison to weight interpolation, we scale the $\alpha$ parameter such that the scaled $\alpha=0$ and $\alpha=1$ correspond to the models with attribute score equal to that of the fine-tuned endpoint models. We evaluate these interpolation methods by their proximity to the ground truth interpolated models, which are Llama-2-7b models fine-tuned with $\alpha$ fraction data from class $1$ and $1-\alpha$ fraction data from class 0.
@@ 3 Continuous Language Model Interpolation
We begin by investigating the linear interpolations between each pair of fine-tuned anchor models (3.1). We then extend this analysis to the convex hull of anchor models for multiple attributes (3.2).
Figure 2: Interpolated models recover custom fine-tuned models across the interpolation range. We show the attribute scores for our interpolation framework with weight $\alpha$ compared to DExperts (Liu et al., 2021) and model arithmetic (Dekoninck et al., 2024) with $\alpha$ scaled such that the scaled $\alpha=0$ and $\alpha=1$ models have the same score as the fine-tuned endpoint models. Weight interpolation most closely follows the trend of the ground truth fine-tuned models. We find similar results for other language models in Appendix A.7.
Table 1: Weight interpolation best approximates custom fine-tuned models while producing high-quality text. For every attribute, we report the mean absolute error (MAE) between the attribute scores of the custom fine-tuned models and the models for each approach with corresponding $\alpha$ attribute parameters. We evaluate the text quality using WikiText perplexity and n-gram diversity scores. Weight interpolation produces text significantly closer in attribute score to the fine-tuned models with similar perplexity and better diversity than previous approaches. Appendix A.7 shows results for other language models and and we report additional diversity metrics in Table 6.
Attribute score mean absolute error (MAE)
WikiText
Perplexity
Diversity
Sentiment Politeness Formality Simplicity Humor Average Dist-1 Dist-2 Dist-3
Fine-tuned (ground truth) - - - - - - 5.040 0.930 0.941 0.896
DExperts (Liu et al., 2021) 0.066 0.132 0.105 0.161 0.080 0.109 5.031 0.930 0.892 0.835
Model arithmetic (Dekoninck et al., 2024) 0.075 0.177 0.088 0.143 0.276 0.152 4.900 0.923 0.936 0.889
Weight interpolation 0.036 0.041 0.006 0.066 0.105 0.051 4.947 0.934 0.939 0.891